# Creation of ultracold heteronuclear -wave Feshbach molecules

### v2

@@ Abstract
Currently, the most studied ultracold atomic systems with higher partial wave interactions is the single-component fermions, which lack -wave interactions but can exhibit tunable -wave interactions near -wave FRs. A range of interesting topics has been explored with such systems, including the association of -wave FMs [11, 22], the -wave contact [44, 28], ultracold fermions with unitary -wave interactions [39], and emergent -wave interactions in low dimensions [23]. Another notable study involving high partial wave interactions is the recent successful creation of a Bose-Einstein condensate (BEC) of -wave FMs starting from a Cs atomic BEC in a two-dimensional trap [2].
@@ Acknowledgements.
We are grateful to Bin Zhu, Yue Cui, Zhendong Zhang, and Shizhong Zhang for valuable discussions, aw well as Xinyuan Gao and Yangqian Yan for their careful reading of our manuscript. This work is supported by Quantum Science and Technology–National Science and Technology Major Project of China (2024ZD0300600), the Hong Kong RGC General Research Fund (Grants 14304323 and 14302722) and the Collaborative Research Fund (Grant No. C4050-23G), and Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2303002).References
- [1] Note: See supplementary material, which includes Refs. [32,34]. Cited by: Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.

### v3

@@ Abstract
Currently, the most studied ultracold atomic systems with higher partial wave interactions is the single-component fermions, which lack -wave interactions but can exhibit tunable -wave interactions near -wave FRs. A range of interesting topics has been explored with such systems, including the association of -wave FMs [11, 24], the -wave contact [46, 19, 49, 20, 30] ultracold fermions with unitary -wave interactions [41], and emergent -wave interactions in low dimensions [25]. Another notable study involving high partial wave interactions is the recent successful creation of a Bose-Einstein condensate (BEC) of -wave FMs starting from a Cs atomic BEC in a two-dimensional trap [2].
@@ Acknowledgements.
We are grateful to Bin Zhu, Yue Cui, Zhendong Zhang, Mingyang Liu and Shizhong Zhang for valuable discussions, as well as Xinyuan Gao and Yangqian Yan for their careful reading of our manuscript. This work is supported by Quantum Science and Technology–National Science and Technology Major Project of China (2024ZD0300600), the Hong Kong RGC General Research Fund (Grants 14304323 and 14302722) and the Collaborative Research Fund (Grant No. C4050-23G), and Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2303002).References
- [1] Note: See supplementary material. Cited by: Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules, Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [19] (2016-01) Concept of a contact spectrum and its applications in atomic quantum hall states. Phys. Rev. Lett. 116, pp. 045301. External Links: Document, Link Cited by: Creation of ultracold heteronuclear -wave Feshbach molecules.
- [20] (2021-10) -Wave contacts of quantum gases in quasi-one-dimensional and quasi-two-dimensional traps. Phys. Rev. A 104, pp. 043303. External Links: Document, Link Cited by: Creation of ultracold heteronuclear -wave Feshbach molecules.
@@
- [49] (2017-06) Contact matrix in dilute quantum systems. Phys. Rev. A 95, pp. 062702. External Links: Document, Link Cited by: Creation of ultracold heteronuclear -wave Feshbach molecules.